Maggie Motors, Inc. (MMI) currently has a debt-to-equity ratio of 0.4 and a total market value of $300 million. MMI believes that increasing leverage (by taking on debt and repurchasing an equal dollar amount of shares) will increase firm value; consequently, TTI has set a target debt-to-equity ratio of 0.6, which they plan to maintain permanently. If MMI’s marginal tax rate is 25%, then present value of the tax shield it will create?
Answer #1

If the total debt to equity ratio is .4

It will mean that total debt to Capital ratio would be= .4/(.4+1)= (.4/1.4)=.2857

It is given that total market value is 300 million.

When the debt to equity ratio .6

It will mean that total debt to Capital ratio

= .6+(1+.6)= (.6/1.6)=.375

Total value of company= 300*.375/.2857= 393,770

Total value of debt= .375*393770= 147664

Total value of interest tax shield= (value of the debt x Tax rate)

= (147664*25%)=36916.